# Boundary Attention: Learning to Localize Boundaries under High Noise

### v2

@@ Abstract
We present a differentiable model that infers explicit boundaries, including curves, corners and junctions, using a mechanism that we call boundary attention. Boundary attention is a boundary-aware local attention operation that, when applied densely and repeatedly, progressively refines a field of variables that specify an unrasterized description of the local boundary structure in every overlapping patch within an image. It operates in a bottom-up fashion, similar to classical methods for sub-pixel edge localization and edge-linking, but with a higher-dimensional description of local boundary structure, a notion of spatial consistency that is learned instead of designed, and a sequence of operations that is end-to-end differentiable. We train our model using simple synthetic data and then evaluate it using photographs that were captured under low-light conditions with variable amounts of noise. We find that our method generalizes to natural images corrupted by real sensor noise, and predicts consistent boundaries under increasingly noisy conditions where other state-of-the-art methods fail.
@@ 1 Introduction
Earlier work by Canny [3] and many others [5, 16, 10, 9, 15] explored the detection of parametric edge models, and there are a variety of bottom-up algorithms that try to connect them by encouraging geometric consistency via edge-linking or message-passing. But since the dawn of deep learning, boundaries have almost exclusively been represented using discrete, rasterized maps, and spatial-consistency mechanisms that were previously based on explicit curve geometry have largely been replaced by black-box neural networks.
In this paper, we take inspiration from early computer vision work and revisit the task of finding unrasterized boundaries via bottom-up geometric processing. But unlike early work, we want to leverage the speed and power of neural networks to learn these processes instead of hand-crafting them.
We focus our attention on the low-level task of finding boundaries that separate regions of uniform color, as depicted by the toy problem in Figure 1. This task becomes difficult at high noise levels, especially within local patches, and we address it by creating a model that can learn to exploit a wide range of low-level cues, such as curves being predominantly smooth with low-curvature; curves tending to meet at corner and junction points; contours tending to have consistent contrast polarity throughout their extent; colors tending to vary smoothly at locations away from boundaries; and so on.
@@
Figure 1: Process overview. The image unfolds into stride-1 patches, and boundary attention operates iteratively on their embeddings to produce for each patch: (i) a parametric three-way partitioning, and (ii) a parametric windowing function that defines its effective patch size. (Figure 2 shows parameterization details.) This output field implies a variety of global maps, shown in clockwise order: a boundary-aware smoothing of the input colors; an unsigned boundary-distance map; a boundary map; and a map of spatial affinities between any query point and its neighbors.
We design our model so that all of its components are local and invariant to discrete spatial shifts, enabling it to be trained on small-sized images and then deployed on much larger and differently-shaped ones. By tying the weights across multiple portions of the network, we increase its resilience to noise which we validate with ablations. Our resulting model is very compact, comprising only 207k parameters, and it runs several times faster than many alternatives. Further, our model can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise. Despite this simplicity, we find that our model can generalize to captured photographs and novel noise processes surprisingly well. We evaluate our model on real noisy images and show that it can capture fine details that other methods miss, and moreover maintain its predictions over high levels of noise.
@@
    We propose a novel network design that explicitly models boundaries using a boundary attention mechanism that can be used in any deep learning framework.
@@
    We demonstrate that our model based on boundary attention is more effective than current state-of-the-art methods at precisely and consistently finding boundaries at high levels of sensor noise (Figure 6).
Figure 2: Parameterization details. Left: Each patch $k$ is associated with an unrasterized three-way partitioning of its area (colored blue, orange and purple here). The partitioning parameters comprise a vertex $(u,v)$, orientation $\theta$, and angles $(\omega_{1},\omega_{2},\omega_{3})$, defined up to scale. A: A walk through junction space by linearly interpolating between junctions is spatially smooth, and can represent edges, bars, corners, Y-junctions, T-junctions and uniform regions. B: Each junction is modulated through a learned windowing function. The windowing parameters $\mathbf{p}=(p_{1},p_{2},p_{3})$ are convex weights over a dictionary of binary pillboxes. See the supplement for an interactive visualization.
@@ 2 Related Work
Recently, the field of junctions (FoJ) [25] introduced a way to detect more-complicated local structures while also representing them parametrically, without rasterization. By using its unrasterized representations in a customized non-convex optimization routine, the method can detect curves, corners and junctions with high precision and with unprecedented resilience to noise. Our model is strongly inspired by this work, and it improves upon it by introducing an enhanced family of parameterized geometric primitives (Figure 2) and by replacing its hand-designed objective and associated non-convex optimization routine with a learned model that is fast, feed-forward and differentiable. We also tackle the challenge of representing images that vary in scale across different areas. Unlike many low-level methods that assess image edges at a single, frequently manually chosen scale, our approach considers variations in scale throughout the image. The field of junctions, for example, requires setting a single patch size that junctions are optimized over, where larger patch sizes allow for better resilience to noise at the expense of memory and detail preservation. In contrast we equip each local region with a function to modulate its spatial extent.
Our work is very different from the recent trend in computer vision of using deep networks to detect boundaries that are semantically meaningful. Examples include HED [27], EDTER [19], and UAED [28].
These models are trained using annotated segmentation datasets like BSDS500 [2] and NYUDv2 [22], and they aim to identify boundaries between semantic objects (e.g., people, animals) while ignoring boundaries that are inside of these objects (e.g., stripes on a shirt). Their large capacities and their end-to-end training allow them to exploit not only low-level cues but also high-level cues such as object and scene familiarity.
In contrast to this recent trend, we follow earlier work by focusing entirely on identifying boundaries from the bottom up, using low-level cues alone. Since our model is trained using simple synthetic data, it does not exploit object familiarity or learn to ignore intra-object edges. We find that this approach maintains some advantages: instead of producing a pixelized boundary map, our model produces a rich field of unrasterized primitives that provide better precision as well as information about local grouping. Moreover, despite not emphasizing top-down information, we find it performs better than deep models in the presence of high noise levels. This suggests that future work may benefit from incorporating some of our architectural elements into larger and higher-capacity systems, in particular to improve interpretability, efficiency, generalization and task-transfer.
@@ 3 Representation
Our system is depicted in Figure 1. It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with stride-1 patch embeddings. Critically, each embedding explicitly encodes a tuple of values that specifies the geometric structure and spatial extent of the local boundaries within a patch. Because the embeddings are forced through this geometric bottleneck, we refer to our model’s core mechanism as boundary attention.
We instantiate our bottleneck using a learned linear mapping from the embedding dimension to a predefined lower-dimensional space of unrasterized boundary patterns that we call junction space. Junction space has the benefit of specifying local boundary patterns without rasterization and thus with unlimited spatial precision. As depicted in Figure 2 and described in [25], it also has the benefit of including a large family of local boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions. We also modulate the spatial extent of each junction by learning an associated windowing function, described in Section 3.1.
To enable communication between patch embeddings, each of which corresponds to a local region of an image, we leverage the idea that each region of the image is described by multiple patches, and that these patches should agree on compatible representations of the global image structure. The ideal embedding for each patch should aptly represent its local geometry while agreeing with the structural claims of neighboring patches.
The next section introduces our parameterization of junction space and some associated operators. Then Section 4 describes the architecture that we use to transform an image into its junction space representation.
@@ 3.1 Boundary Primitives
Throughout this paper we use parentheses $(x)$ for continuous signals defined on the 2D image plane $[0,W]\times[0,H]$ and square brackets $[n]$ for discrete signals defined on the pixel grid. We use $c[n]$ for the coordinates of the $n$th pixel’s center.
Denote the $K$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{R}^{K}$ is the vector image value at the discrete pixel grid index $n$. Our approach is to analyze the image into a field of dense, stride-1 overlapping local patches, each having a square support $\Omega_{k}(x)$ for the $k$th patch centered at the $k$th pixel.
There are many ways to partition a local region $\Omega_{k}(x)$, and one can define parametric families of partitions. For example the set of oriented lines provides a two-parameter family of partitions, with each member of the family separating the region into points that lie on one side of a line or the other. This family of partitions would be appropriate for describing edges. Here we define a larger family of partitions that encompasses a greater variety of local boundary structures.
As depicted in the right of Figure 2, our partitions are parameterized by $\mathbf{g}\in\mathbb{R}^{2}\times\mathbb{S}^{1}\times\triangle^{2}$, where $\mathbb{S}^{1}$ is the unit circle and $\triangle^{2}$ is the standard $2$-simplex. We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\omega_{3})$ are barycentric coordinates (defined up to scale) for the three relative angles, ordered clockwise starting from $\theta$. Our convention is to express the vertex coordinates relative to the center of region $\Omega_{k}(x)$, which is located at $c[k]$, and we denote that the vertex is free to move outside of this region. We also note that up to two angles $\omega_{j}$ can be zero. This all makes it possible to represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions), and do so with unlimited spatial resolution.
@@
In order to enable the size of each region $\Omega_{n}$ to be able to adapt to the local geometry and noise conditions, we equip each one with a parameterized local windowing function $w_{k}(x;\mathbf{p})\in[0,1]$, with parameters $\mathbf{p}\in{\cal P}=\triangle^{W-1}$ that are the coefficients of a convex combination of $W$ square pillbox functions. That is,
@@
[…]
Figure 3: Example of our model’s output, with examples from two different regions. Top row: Some of each region’s overlapping input patches, their corresponding outputs (visualized in the style of Figure 2), and three types of per-patch attributes that the outputs imply: unsigned distance; boundaries; and gathered wedge colors. Bottom row: Four types of global maps that are implied by accumulating values from the output field and rendered patches. See the supplement for interactive visualizations.
@@ 3.2 Gather and Slice Operators
Note that only patches that contain $c[n]$ contribute to the sum, i.e., the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$. The expressions for the pixel-wise distance map variance $\nu_{d}[n]$, and for the feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are defined accordingly and included in the supplement. Intuitively, slicing $\mathbf{f}_{kj}$ can be thought of globally accumulating information from each patch’s junction representation.
@@ 3.3 Visualizing Output
At our network’s output, we are provided with a field of tuples representing the junction and window parameters $\{(\mathbf{g}[k],\mathbf{p}[k]\}$ for all stride-1 patches of the input image. Though these outputs are defined continuously, we can visualize them in Figure 3 by rendering the windowing functions $w_{k}(x;\mathbf{p})$ (second column), and the binary junction wedge supports $s_{kj}(x;\mathbf{g})$ colored purple, orange, and blue (third column), along with the input image patches (first column). To the left, we show the input image and the regions from which the patches were extracted from.
Additionally, we can use the learned junction parameters to render unsigned distance patches $d_{k}(x;\mathbf{g})$ (fourth column), boundary patches $b_{k}(x;\mathbf{g})$ (fifth column), and the binary junction wedges $s_{kj}(x;\mathbf{g})$ colored with their respective gathered colors (sixth column). All of the patches are defined continuously, so we can render them at any resolution.
@@
The output contains much more information than these fields of means. For any query pixel $n$, we can probe the wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ that contain it, thereby obtaining a spatial attention map $a_{k}(x)$ that surrounds the query pixel. This is the boundary-aware spatial kernel that turns a neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the value $\bar{\mathbf{f}}[n]$:
@@
The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6, and its maximum diameter is twice that of $\Omega(x)$. The spatial affinity maps for two probed points are shown in the leftmost image of the bottom row of Figure 3. Like slicing, probing the affinity map around a pixel is a form of global accumulation that takes values defined per-patch and synthesizes them into a global representation.
Figure 4: Model Architecture. All blocks are invariant to discrete spatial shifts, and only colored blocks are learned. Orange blocks operate at individual locations $n$, while blue ones operate on small spatial neighborhoods. Symbol $\oplus$ is concatenation, and gather and slice operators (Eqs. 46) are depicted at right. The first iteration uses $\boldsymbol{\gamma}^{0}[n]=\boldsymbol{\gamma}_{0}[n]$, $\bar{\mathbf{f}}^{0}[n]=\mathbf{f}[n]$, and $\boldsymbol{\pi}^{0}[n]=\boldsymbol{\pi}_{o}$ with $\boldsymbol{\pi}_{o}$ learned across the training set. Boundary attention repeats $T=8$ times, with one set of weights for the first four iterations and another for the last four iterations. There are $207$k trainable parameters in total.

### v3

# Boundary Attention: Learning curves, corners, junctions and grouping

@@ Abstract
We present a lightweight network that infers grouping and boundaries, including curves, corners and junctions. It operates in a bottom-up fashion, analogous to classical methods for sub-pixel edge localization and edge-linking, but with a higher-dimensional representation of local boundary structure, and notions of local scale and spatial consistency that are learned instead of designed. Our network uses a mechanism that we call boundary attention: a geometry-aware local attention operation that, when applied densely and repeatedly, progressively refines a pixel-resolution field of variables that specify the boundary structure in every overlapping patch within an image. Unlike many edge detectors that produce rasterized binary edge maps, our model provides a rich, unrasterized representation of the geometric structure in every local region. We find that its intentional geometric bias allows it to be trained on simple synthetic shapes and then generalize to extracting boundaries from noisy low-light photographs.
@@ 1 Introduction
Earlier work by Canny [2] and many others [4, 15, 9, 8, 14] explored the detection of unrasterized parametric edge models, and there are a variety of bottom-up algorithms that try to connect them by encouraging geometric consistency via edge-linking or message-passing. But since the dawn of deep learning, boundaries have almost exclusively been represented using discrete, rasterized maps; and spatial-consistency mechanisms that were previously based on explicit curve geometry have largely been replaced by black-box neural networks.
In this paper, we take inspiration from early computer vision work and revisit the task of finding unrasterized boundaries via bottom-up geometric processing. But unlike early work, we leverage self-attention to learn these processes instead of hand-crafting them, thereby combining the benefits of geometric modeling with the efficiency and representational power of deep learning.
We focus on the low-level task of finding boundaries that separate regions of uniform color, as depicted by the toy problem in Figure 1. This task becomes difficult at high noise levels, especially within local patches, and we address it by creating a model that can learn to exploit a wide range of low-level cues, such as curves being predominantly smooth with low-curvature; curves tending to meet at corner and junction points; contours tending to have consistent contrast polarity throughout their extent; and colors tending to vary smoothly at locations away from boundaries.
@@
Figure 1: Pipeline overview. The image unfolds into stride-1 patches, and boundary attention operates iteratively on their embeddings to produce for each patch: (i) a parametric three-way partitioning, and (ii) a parametric windowing function that defines its effective patch size. (Figure 2 shows parameterization details.) This output field implies a variety of global maps, shown in clockwise order: a boundary-aware smoothing of the input colors; an unsigned boundary-distance map; a boundary map; and a map of spatial affinities between any query point and its neighbors.
We intentionally design our model with components that are local and invariant to discrete spatial shifts, enabling it to be trained on small-sized images and then deployed on much larger and differently-shaped ones. By tying the weights across multiple portions of the network, we increase its resilience to noise, which we validate through ablations. Our resulting model is very compact, comprising only 207k parameters, and it runs much faster than comparable optimization-based approaches. Further, our model can be trained to a useful state with very simple synthetic data, made up of random circles and triangles that are uniformly colored and then corrupted by noise. Despite the simplicity of this training data, we find that the model’s learned internal activations exhibit intuitive behavior, and that the model generalizes to real world photographs, including those taken in low-light and have substantial shot noise.
@@
    We introduce a bottom-up, feedforward network that decomposes an image into an underlying field of local geometric primitives that explicitly identify curves, corners, junctions and local grouping.
@@
    We do this by introducing a new parameterization for local geometric primitives, and a new self-attention mechanism that we call boundary attention.
  3. 3.
    We identify the best architecture among a family of candidates, and we show that it generalizes from simple synthetic images to real low-light photographs.
Figure 2: Parameterization details. Left: Each patch $k$ is associated with an unrasterized three-way partitioning of its area (colored blue, orange and purple here). The partitioning parameters comprise a vertex $(u,v)$, orientation $\theta$, and angles $(\omega_{1},\omega_{2},\omega_{3})$, defined up to scale. A: A walk through junction space by linearly interpolating between junctions is spatially smooth, and can represent edges, bars, corners, Y-junctions, T-junctions and uniform regions. B: Each junction is modulated through a learned windowing function. The windowing parameters $\mathbf{p}=(p_{1},p_{2},p_{3})$ are convex weights over a dictionary of binary pillboxes.
@@ 2 Related Work
Recently, the field of junctions (FoJ) [23] introduced a way to detect more-complicated local structures while also representing them parametrically, without rasterization. By using its unrasterized representations in a customized non-convex optimization routine, the method can detect curves, corners and junctions with high precision and with unprecedented resilience to noise. Our model is strongly inspired by this work, and it improves upon it by introducing an enhanced family of parameterized geometric primitives (Figure 2) and by replacing its hand-designed objective and associated non-convex optimization routine with a learned model that is fast, feed-forward and differentiable. We also tackle the challenge of representing images that vary in scale across different areas: whereas FoJ requires setting a single patch size for the entire image, our model learns to choose patch sizes adaptively in response to the input image.
We emphasize that our work is very different from a recent trend in computer vision of using deep networks to detect boundaries that are semantically meaningful (e.g.[21, 18, 27]). These models aim to identify boundaries between semantic concepts, such as people and animals, while ignoring boundaries that are inside of these concepts, such as stripes on a shirt. In contrast to this trend, we follow earlier work by focusing entirely on identifying boundaries from the bottom up, using low-level cues alone. Since our model is trained using simple synthetic data, it does not exploit object familiarity or learn to ignore intra-object contrast. This approach has advantages: It is not specialized to any predetermined set of semantic concepts or tasks, and instead of producing a pixelized boundary map, it produces a field of unrasterized primitives that provide better precision as well as richer information about local grouping. We leave for future work the exploration of how to combine our bottom-up model with other cues that are top-down and semantic.
@@ 3 Representation
Our system is depicted in Figure 1. It uses neighborhood cross-attention, a patch-wise variant of cross-attention, with $D$-dimensional, stride-1 embeddings. Critically, each $D$-dimensional embedding explicitly encodes a tuple of values that specifies the geometric structure and spatial extent of the unrasterized local boundaries within a patch.
Our model relies on a learned (linear) mapping from the embedding dimension $D$ to a hand-crafted, lower-dimensional space of unrasterized boundary patterns that we call junction space. Junction space has the benefit of specifying per-patch boundary patterns without rasterization and thus with unlimited spatial precision. As depicted in Figure 2 and described in [23], it also has the benefit of including a large family of per-patch boundary patterns, including uniformity (i.e., absence of boundaries), edges, bars, corners, T-junctions and Y-junctions. Different from [23], we additionally modulate the spatial extent of each patch by learning an associated windowing function, to be described in Section 3.1.
To enable communication between patch embeddings, each of which corresponds to a local patch, we leverage the idea that each image point is covered by multiple patches, and that overlapping patches must agree in their regions of overlap. This has the effect of tying neighboring patches together, analogous to cliques in a Markov random field.
We refer to the core mechanism of our model as boundary attention. We introduce it by first defining our hand-crafted parameterization of junction space and some associated operators. Then Section 4 describes the architecture that we use to transform an image into a coherent junction-space representation.
@@ 3.1 Boundary Primitives
We use parentheses $(x)$ for continuous signals defined on the 2D image plane $[0,W]\times[0,H]$ and square brackets $[n]$ for discrete signals defined on the pixel grid. We use $c[n]$ for the coordinates of the pixel with integer index $n$.
Denote the $C$-channel input image by $\{\mathbf{f}[n]\}$, where $\mathbf{f}[n]\in\mathbb{Q}^{C}$ is the vector image value at the discrete pixel grid index $n$. Our approach is to treat the image as a field of dense, stride-1 overlapping local patches. We use $\Omega_{k}(x)$ to denote the spatial support of the patch that is centered at the pixel whose integer index is $k$.
There are many ways to partition a local patch $\Omega_{k}(x)$, and one can define parametric families of partitions. For example the set of oriented lines provides a two-parameter family of partitions, with each member of the family separating the region into points that lie on one side of a line or the other. This family of partitions would be appropriate for describing edges. Here we define a larger family of partitions that encompasses a greater variety of local boundary structures.
As depicted in the right of Figure 2, our partitions are parameterized by $\mathbf{g}\in\mathbb{R}^{2}\times\mathbb{S}^{1}\times\triangle^{2}$, where $\mathbb{S}^{1}$ is the unit circle and $\triangle^{2}$ is the standard $2$-simplex. We use the notation $\mathbf{g}=(\boldsymbol{u},\theta,\boldsymbol{\omega})$, where $\boldsymbol{u}=(u,v)\in\mathbb{R}^{2}$ is the vertex, $\theta\in\mathbb{S}^{1}$ is the orientation, and $\boldsymbol{\omega}=(\omega_{1},\omega_{2},\omega_{3})$ are barycentric coordinates (defined up to scale) for the three relative angles, ordered clockwise starting from $\theta$. Our convention is to express the vertex coordinates relative to the center of region $\Omega_{k}(x)$, which is located at $c[k]$, and we note that the vertex is free to move outside of this region. We also note that up to two angles $\omega_{j}$ can be zero. This all makes it possible to smoothly represent a variety of partition types, including edges, bars, corners, 3-junctions and uniformity (i.e., trivial or singleton partitions), and do so with unlimited spatial resolution.
@@
In order to enable the size of each region $\Omega_{k}$ to adapt to the local geometry and noise conditions, we equip each one with a parameterized local windowing function $w_{k}(x;\mathbf{p})\in[0,1]$, with parameters $\mathbf{p}\in{\cal P}=\triangle^{W-1}$ that are the coefficients of a convex combination of $W$ square pillbox functions. That is,
@@
where $\|\cdot\|_{\infty}$ is the $\ell^{\infty}$-norm, and $\mathbf{1}[\cdot]$ is the indicator function that returns $1$ if the argument is true; and 00 otherwise. In our experiments we use $W=3$ and diameters $\mathbf{D}=(3,9,17)$. Figure 3 shows some examples.
Figure 3: Example of our model’s output, with examples from two different regions. Top row: Some of each region’s overlapping input patches, their corresponding outputs (visualized in the style of Figure 2), and three types of per-patch attributes that the outputs imply: unsigned distance; boundaries; and gathered wedge colors. Bottom row: Four types of global maps that are implied by accumulating values from the output field and rendered patches.
@@ 3.2 Gather and Slice Operators
Here, the sums are over $\{k\mid\Omega_{k}\ni c[n]\}$ so that only patches that contain $c[n]$ contribute to the sum. Similar expressions for pixel-wise distance map variance $\nu_{d}[n]$ and feature variance $\nu_{f}[n]$, which is computed across patches containing $n$ and across their $K$ channels, are included in the supplement. Intuitively, slicing represents an accumulation of regional information into a pixel, as dictated by the partitions of all of the patches that contain the pixel.
@@ 3.3 Visualizing Output
Our network’s output is a field of tuples representing the junction and windowing parameters $\{(\mathbf{g}[k],\mathbf{p}[k]\}$ for all stride-1 patches of the input image. We visualize them in Figure 3 by rasterizing the continuous windowing functions $w_{k}(x;\mathbf{p})$ (second column) and binary wedge supports $s_{kj}(x;\mathbf{g})$, which are colored purple, orange, and blue (third column). To the left, we show the input image and the regions from which the patches were extracted.
Additionally, we can use the output junction parameters to rasterize unsigned distance patches $d_{k}(x;\mathbf{g})$ (fourth column), boundary patches $b_{k}(x;\mathbf{g})$ (fifth column), and wedge supports $s_{kj}(x;\mathbf{g})$ that are re-colored with their respective wedge features $\mathbf{f}_{kj}$ (sixth column). Note that all of these are defined continuously, so can be rasterized at any resolution.
@@
For any query pixel $k$, we can also probe the containing wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ to compute a spatial affinity map $a_{k}(x)$ that surrounds the query pixel. This represents the affinity between point $c[k]$ and a surrounding neighborhood with diameter that is twice that of $\Omega(x)$. It is also the boundary-aware spatial kernel that turns the neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the gathered value $\bar{\mathbf{f}}[n]$ via
@@
The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6. The spatial affinity maps for two probed points are shown in the leftmost image of the bottom row of Figure 3. Like slicing, querying the affinity map at a pixel is a form of regional accumulation from patches to a pixel.
Figure 4: Model Architecture. All blocks are invariant to discrete spatial shifts, and only colored blocks are learned. Orange blocks operate at individual locations $n$, while blue ones operate on small spatial neighborhoods. Symbol $\oplus$ is concatenation, and gather and slice operators (Eqs. 46) are depicted at right. The first iteration uses $\boldsymbol{\gamma}^{0}[n]=\boldsymbol{\gamma}_{0}[n]$, $\bar{\mathbf{f}}^{0}[n]=\mathbf{f}[n]$, and $\boldsymbol{\pi}^{0}[n]=\boldsymbol{\pi}_{o}$ with $\boldsymbol{\pi}_{o}$ learned across the training set. Boundary attention repeats $T=8$ times, with one set of weights for the first four iterations and a separate set of weights for the last four iterations, resulting in 207k trainable parameters total.